# A Survey of Foundation Models for IoT: Taxonomy and Criteria-Based Analysis

### v1

@@ 6. Criterion IV: Security and Privacy
In this section, we summarize approaches to enhancing IoT security and privacy with foundation models in two contexts: (1) using foundation models to protect the security and privacy of IoT systems without directly participating in downstream tasks (e.g., human activity recognition, robot control), and (2) securing FM-based IoT systems where the foundation model plays a central role in downstream tasks.
@@
Privacy Considerations for FM-based IoT system. Two techniques are commonly used by the literature to protect privacy in FM-based IoT systems: (1) Using less interpretable data modalities: Selecting data types that are difficult for humans to interpret, such as vibration data instead of video for tasks like human activity monitoring, can improve privacy (Yang et al., 2025). However, this approach requires foundation models capable of understanding such non-human-interpretable data, which is more challenging due to the limited availability of labeled training datasets. (2) Data obfuscation in edge–cloud architectures: When the edge-cloud collaboration system (Yao et al., 2022; Tian et al., 2024) is leveraged, data should be obfuscated on local devices, using rule-based methods, machine learning models, or lightweight foundation models, before being transmitted to the cloud (Ouyang and Srivastava, 2024; Benazir and Lin, 2024; Yuan et al., 2024). The cloud hosts more powerful FMs for task-specific reasoning, but transmitting raw data increases exposure to privacy risks during transit and in cloud environments, where broader access makes data more vulnerable to attacks.
@@ 7. Evaluation
We also summarize specific evaluation metrics from the reviewed papers for each performance criterion discussed in this work: (1) Efficiency: Convergence speed (e.g., number of epochs to convergence) assesses the training efficiency of fine-tuning-based methods (Kimura et al., 2024a, b). Inference latency measures the time required for the FM-based IoT system to complete a downstream task during inference (Yang et al., 2023a). Memory usage measures the amount of memory consumed during inference (Worae et al., 2024). (2) Context-Awareness: Executability rate evaluates how likely the model’s response can be executed in a specific environment (Liu et al., 2024b). Personalization score assesses how well the response is tailored to an individual’s unique characteristics (e.g., persona, background, schedule) (Yang et al., 2025). (3) Safety: Safety rate quantifies the likelihood that a model’s response violates predefined safety constraints (Ma et al., 2025). (4) Security and Privacy: For security, the common task is anomaly detection and the corresponding metrics are classification based metrics (e.g., F1, precision, recall) (Li et al., 2024f; Houssel et al., 2024; Diaf et al., 2024; Almodovar et al., 2022; Aghaei et al., 2022). For privacy, evaluation typically involves the attack accuracy, which measures how accurately an adversarial model can recover sensitive information that is intended to be protected (Yang and Ardakanian, 2023; Wang et al., 2024a).

### v2

@@ 7. Criterion IV: Security and Privacy
In this section, we summarize approaches to enhancing IoT security and privacy with foundation models in two contexts: (1) using foundation models to protect the security and privacy of IoT systems without directly participating in downstream tasks (e.g., human activity recognition, robot control), and (2) securing FM-based IoT systems where the foundation model plays a central role in downstream tasks. Table 6 provides also an overview of these methods, including their descriptions and the paradigms (prompt-based, agent-based, training-based) to which they are mostly commonly applied.
Table 6. Summary of Security&Privacy-Enhancing Methods, with Descriptions and Typically Applicable Paradigms (Prompt-Based, Agent-Based, Training-Based).
Typically Applicable Paradigm
Method Description Prompt Agent Training
FMs for Anomaly Detection
Detecting potential anomalies, such as unusual
activities in system or network logs and malicious
sensor data-sharing requests.
FMs for Data Obfuscation
Transforming sensitive data into a less meaningful
or recognizable form to prevent unauthorized access
and ensure compliance with data protection regulations.
@@
Privacy Considerations for FM-based IoT system. Two techniques are commonly used by the literature to protect privacy in FM-based IoT systems: (1) Using less interpretable data modalities: Selecting data types that are difficult for humans to interpret, such as vibration data instead of video for tasks like human activity monitoring, can improve privacy (Yang et al., 2025a). However, this approach requires foundation models capable of understanding such non-human-interpretable data, which is more challenging due to the limited availability of labeled training datasets. (2) Data obfuscation in edge–cloud architectures: When the edge-cloud collaboration system (Yao et al., 2022; Tian et al., 2024) is leveraged, data should be obfuscated on local devices, using rule-based methods, machine learning models, or lightweight foundation models, before being transmitted to the cloud (Ouyang and Srivastava, 2024; Benazir and Lin, 2024; Yuan et al., 2024). The cloud hosts more powerful FMs for task-specific reasoning, but transmitting raw data increases exposure to privacy risks during transit and in cloud environments, where broader access makes data more vulnerable to attacks.
@@ 8. Evaluation
We also summarize specific evaluation metrics from the reviewed papers for each performance criterion discussed in this work: (1) Efficiency: Convergence speed (e.g., number of epochs to convergence) assesses the training efficiency of fine-tuning-based methods (Kimura et al., 2024a, b). Inference latency measures the time required for the FM-based IoT system to complete a downstream task during inference (Yang et al., 2023a). Memory usage measures the amount of memory consumed during inference (Worae et al., 2024). (2) Context-Awareness: Executability rate evaluates how likely the model’s response can be executed in a specific environment (Liu et al., 2024b). Personalization score assesses how well the response is tailored to an individual’s unique characteristics (e.g., persona, background, schedule) (Yang et al., 2025a). (3) Safety: Safety rate quantifies the likelihood that a model’s response violates predefined safety constraints (Ma et al., 2025). (4) Security and Privacy: For security, the common task is anomaly detection and the corresponding metrics are classification based metrics (e.g., F1, precision, recall) (Li et al., 2024f; Houssel et al., 2024; Diaf et al., 2024; Almodovar et al., 2022; Aghaei et al., 2022). For privacy, evaluation typically involves the attack accuracy, which measures how accurately an adversarial model can recover sensitive information that is intended to be protected (Yang and Ardakanian, 2023; Wang et al., 2024a).